# Sample, estimate, aggregate: A recipe for causal discovery foundation models

### v2

@@ 3.4 Theoretical interpretation
Marginal graph resolution
It is well-established that estimates of causal graphs over subsets of variables can be “merged” into consistent graphs over their union (Faller et al., 2023), and various algorithms have been proposed towards this task (Tillman et al., 2008; Huang et al., 2020a). Our main theoretical contribution is demonstrating that our axial attention model can implement such an algorithm under realistic conditions (description of algorithm and all proofs in Appendix A).
@@
The second consideration is that checking a combinatorial number of subsets is wasteful on smaller graphs and infeasible on larger graphs. In fact, if we only leverage marginal estimates, we must check at least $O(N^{2})$ subsets to cover each edge at least once. To this end, the classical Independence Graph algorithm (Spirtes et al., 2001) motivates statistics such as inverse covariance to initialize the undirected skeleton and reduce the number of independence tests required. This allows us to use marginal estimates more efficiently, towards answering orientation questions.
On identifiability
The goal of this paper is a principled, yet practical framework for causal discovery. Instead of focusing on the identifiability of any particular setting, we offer the following interpretations of our model’s outputs. When all assumptions are upheld, and infinite data are available, the model has the capacity to infer a sound graph, as far as its (Markov/interventional) equivalence class. Given only observational data, the model will attempt to orient all edges, but the output can be interpreted as a member of the MEC. Given interventional data with sufficient interventions, the model should be able to identify the ground truth (Hauser & Bühlmann, 2012). When data do not match a causal discovery algorithm’s assumptions, its performance is inherently an empirical question. In this case, we show empirically that our model still does well (Section 5.1).
@@ 4 Experimental setup
We pretrained Sea models on diverse synthetic datasets. We evaluate on unseen in-distribution and out-of-distribution synthetic datasets, as well as two real biological datasets (Sachs et al., 2005; Replogle et al., 2022), using the versions from Wang et al. (2017); Chevalley et al. (2022). We also include experiments on simulated mRNA datasets in Appendix C.3 (Dibaeinia & Sinha, 2020). All baselines were trained and/or run from scratch on each testing dataset using their published code and hyperparameters, except Avici (their recommended checkpoint was trained on their synthetic train and test sets after publication, Appendix B.2).
4.1 Synthetic data generation
The training datasets were constructed by 1) sampling Erdős-Rényi and scale free graphs with $N=10,20,100$ nodes and $E=N,2N,3N,4N$ expected edges; 2) sampling random instantiations of causal mechanisms (Linear, NN with additive/non-additive Gaussian noise); and 3) iteratively sampling observations in topological order (details in Appendix B.1). For each graph, we generated both observational and interventional datasets with $1000N$ points, either all observational or split equally among observational and perfect single-node interventions. We generated 90 training, 5 validation, and 5 testing datasets for each combination (8160 total). For testing, we also sampled out-of-distribution datasets with 1) Sigmoid and Polynomial mechanisms (additive/non-additive Gaussian noise); and 2) Linear with additive non-Gaussian noise.
4.2 Causal discovery metrics
We report both discrete and continuous metrics, as neural networks can be notoriously uncalibrated (Guo et al., 2017), and arbitrary discretization thresholds reflect an incomplete picture (Schaeffer et al., 2023). Following previous work (Lorch et al., 2022), we compute metrics with respect to the ground truth graph. In the observational setting, the “oracle” expected value of each metric will vary depending on the size of the equivalence class (e.g. if multiple graphs are valid, their mean accuracy is $<1$ with respect to a single ground truth). For all continuous metrics, we exclude the diagonal since several baselines manually set it to zero (Brouillard et al., 2020; Lopez et al., 2022).
SHD: Structural Hamming distance is the minimum number of edge edits required to match two graphs (Tsamardinos et al., 2006). Discretization thresholds are as published or default to 0.5.
mAP: Mean average precision computes the area under precision-recall curve per edge and averages over the graph. The random guessing baseline depends on the positive rate.
@@ 4.2 Causal discovery metrics
Since OA is normalized by $\|E\|$, it is invariant to the positive rate. In contrast to orientation F1 (Geffner et al., 2022), it is also invariant to the assignment of forward/reverse edges as 1/0.
Table 1: Causal discovery results on synthetic datasets. Each setting encompasses 5 distinct Erdős-Rényi graphs (mAP stds all within 0.02). $\dagger$ indicates that this setting is o.o.d. * indicates non-parametric bootstrapping. Runtimes based on 1 CPU and 1 V100 GPU. Baseline details in B.2.
$N$ $E$ Model Linear NN non-add. Sigmoid Polynomial Overall
10 10 Dcdi-G $0.74$ $2.8$ $0.89$ 1.0 $0.46$ $5.8$ $0.41$ $8.9$ 13425.4
Dcdi-Dsf $0.82$ $2.0$ $0.50$ $4.2$ $0.38$ $6.3$ $0.29$ $11.2$ 11449.7
DiffAn $0.25$ $14.0$ $0.12$ $21.8$ $0.24$ $12.0$ $0.20$ $15.0$ 163.1
Avici $0.45$ $8.2$ $0.65$ $5.6$ $0.52$ $8.3$ $0.31$ $9.6$ 1.1
VarSort* $0.70$ $6.0$ $0.83$ $4.2$ $0.52$ $7.6$ $0.40$ $9.3$ 0.1
Fci* $0.52$ $10.0$ $0.40$ $9.8$ $0.56$ $9.1$ $0.41$ $10.0$ 5.7
Gies* $0.81$ $3.6$ $0.71$ $4.8$ $0.70$ $5.9$ $0.61$ $7.1$ 33.8
Sea (Fci) $0.98$ 1.0 $0.88$ $3.8$ $0.83$ 3.9 $0.62$ $6.1$ 3.8
Sea (Gies) 0.99 $1.2$ 0.91 $3.2$ 0.85 $4.0$ 0.70 5.8 3.5
20 80 Dcdi-G $0.46$ $44.0$ 0.82 $37.4$ $0.48$ $44.2$ $0.37$ $59.7$ 18929.9
Dcdi-Dsf $0.48$ $41.2$ $0.74$ 28.4 $0.48$ $43.6$ $0.38$ $57.6$ 11399.6
DiffAn $0.21$ $127$ $0.18$ $145$ $0.22$ $117$ $0.18$ $157$ 558.0
Avici $0.34$ $75.6$ $0.49$ $61.2$ $0.34$ $70.6$ $0.30$ $75.5$ 1.9
VarSort* $0.76$ $44.8$ $0.47$ $65.2$ $0.59$ $63.4$ $0.38$ $75.6$ 0.3
Fci* $0.30$ $75.8$ $0.30$ $74.4$ $0.41$ $72.3$ $0.34$ $76.6$ 187.6
Gies* $0.41$ $70.0$ $0.46$ $67.4$ $0.50$ $65.6$ $0.49$ $68.1$ 814.3
Sea (Fci) $0.80$ $39.8$ $0.70$ $52.0$ $0.74$ 42.9 $0.55$ 57.0 6.7
Sea (Gies) 0.89 26.8 $0.73$ $50.6$ 0.76 $45.0$ 0.65 $60.1$ 4.3
100 400 Dcd-Fg $0.05$ $3068$ $0.10$ $3511$ $0.13$ $3602$ $0.12$ $3317$ 2265.2
Avici $0.12$ $392$ $0.23$ $382$ $0.10$ $399$ $0.03$ $402$ 7.8
VarSort* $0.80$ $225$ $0.34$ $355$ $0.51$ $350$ $0.27$ $380$ 6.7
Sea (Fci) $0.84$ $162$ $0.52$ $339$ $0.63$ $247$ $0.34$ 325 18.6
Sea (Gies) 0.91 117 0.61 267 0.69 218 0.38 $328$ 3.7
@@ 4.3 Baselines
We compare against several deep learning and classical baselines. Dcdi (Brouillard et al., 2020) extracts the causal graph as a parameter of a generative model. The G and Dsf variants use Gaussian or deep sigmoidal flow likelihoods, respectively. Dcd-Fg (Lopez et al., 2022) follows Dcdi-G, but factorizes the graph into a product of two low-rank matrices for scalability. DiffAn (Sanchez et al., 2023) uses the trained model’s Hessian to obtain a topological ordering, followed by a classical pruning algorithm. Avici (Lorch et al., 2022) uses an amortized inference approach to estimate $P(G\mid D)$ over a class of data-generating mechanisms via variational inference. Both Dcd-* and Avici were run with full knowledge of intervention targets.
VarSort (Reisach et al., 2021) sorts nodes by marginal variance and sparsely regresses nodes based on their predecessors. This naive baseline is intended to reveal artifacts of synthetic data generation. Fci, Gies run the FCI and GIES algorithms over all nodes. VarSort, Fci, and Gies were run using non-parametric bootstrapping (Friedman et al., 1999), with 100 subsets of 500 examples each.
@@ 5 Results
Our experiments cover two settings.
@@
    Section 5.2 focuses on the case where we do know (or can estimate) the class of causal mechanisms or exogenous noise. We show that adapting our pretrained models with this information (at low to zero additional training cost) leads to substantial improvement and exceeds the best baseline trained from scratch.
We highlight representative results towards each goal. Full results and ablations are in Appendix C, with additional metrics, baselines, hyperparameters, model components, and design choices.
5.1 Robustness and generalization
@@
Table 1 summarizes our controlled experiments on synthetic data. Sea exceeds all baselines in the Linear, additive Gaussian case, which matches the models’ assumptions exactly (causal discovery algorithms and inverse covariance). In the misspecified (NN) or misspecified and out-of-distribution settings (Sigmoid, Polynomial), Sea also attains the best performance in the majority of cases, even though Dcdi and Avici both have access to the raw data. Furthermore, our models outperform VarSort in every single setting, while most baselines are unable to do so consistently. This indicates that our models do not simply overfit to spurious features of the synthetic data generation process.
@@
Table 3 illustrates that we exceed baselines on single cell gene expression data from CausalBench (Chevalley et al., 2022; Replogle et al., 2022). Furthermore, when we increase the subset size to $b=2000$, we achieve very high precision (0.838) over 2834 predicted edges. Sea runs within 5s on this dataset of 162k cells and $N=622$ genes, while the fastest naive baseline takes 5 minutes and the slowest deep learning baseline takes 9 hours (run in parallel on subsets of genes).
@@
5.2 Low-cost adaptation
We illustrate two strategies that allow us to use pretrained Sea models with different implicit assumptions. First, if two causal discovery algorithms share the same output format, they can be used interchangeably for marginal estimation (though there may be a performance drop if the mistakes themselves are out of distribution). On observational, linear non-Gaussian data, replacing the Ges algorithm with Lingam (Shimizu et al., 2006) is beneficial without any other change (Table 5). The same effect can be observed on Polynomial and Sigmoid non-additive data, when running Fci with a polynomial kernel conditional independence test (Zhang et al., 2011) instead of the Fisherz test (Table 5). Second, we may “finetune” the aggregator using low-cost methods like LoRA (Hu et al., 2022). Specifically, we keep the same training set and objective, and change the global statistic (or discovery algorithm). Here, we show that finetuning our models for distance correlation is beneficial in both cases, and the combination of strategies results in the highest performance overall, surpassing the best baseline trained from scratch (Dcdi-Dsf).

### v3

@@ 3.4 Theoretical interpretation
Marginal graph resolution  It is well-established that estimates of causal graphs over subsets of variables can be “merged” into consistent graphs over their union (Faller et al., 2023; Tillman et al., 2008; Huang et al., 2020). In A.1 and A.2, we describe a simple algorithm towards this task, based on the intuition that edges absent from the global graph should be absent from at least one marginal estimate, and that v-structures present in the global graph are present in the marginal estimates. We then prove Theorem 3.4, which states that the axial attention architecture is well-specified as a model class. That is, there exists a setting of its weights that can map marginal estimates into global graphs. Our construction in A.3 follows the same reasoning steps as the simple algorithm in A.2, and it provides realistic bounds on the model size.
@@
On identifiability  The primary goal of this paper is to develop a practical framework for causal discovery, especially when data assumptions are unknown. Instead of focusing on the identifiability of any particular setting, we provide these interpretations of our model’s outputs, and show empirically that our model respects classic identifiability theory (Section 5.3). The model will always output an orientation for all edges, but the graph can be interpreted as one member of an equivalence class. Metrics can be computed with respect to either the ground truth graph (if identifiable) or the inferred equivalence class, e.g. the implied CPDAG. When data do not match the estimation algorithm’s assumptions, performance is inherently an empirical question, and we show empirically that our model still does well (Section 5.1).
@@ 4 Experimental setup
Our experiments aim to address the three desiderata proposed in Section 2.2 – namely, generalization, adaptability, and emergent few-shot behavior. These experiments span both real and synthetic data. Real experiments quantify the practical utility of this framework, while synthetic experiments allow us to probe and compare each design choice in a controlled setting.
4.1 Datasets
We pretrained Sea models on 6,480 synthetic datasets, which constitute approximately 280 million individual observations, each of 10-100 variables.555 3 mechanisms, 3 graph sizes, 4 sparsities, 2 topologies, $1000N$ examples, 90 datasets $\to$ 280,800,000 examples. For a sense of scale, single cell foundation models are trained on 300K (Rosen et al., 2024) to 30M cells (Cui et al., 2024). To assess generalization and robustness, we evaluate on unseen in-distribution and out-of-distribution synthetic datasets, as well as two real biological datasets (Sachs et al., 2005; Replogle et al., 2022), using the versions from Wang et al. (2017); Chevalley et al. (2025). To probe for emergent few-shot behavior, we down-sample both the training and testing sets. We also include experiments on simulated mRNA datasets with unseen datasets in Appendix C.3 (Dibaeinia & Sinha, 2020).
The training datasets were constructed by 1) sampling Erdős-Rényi and scale free graphs with $N=10,20,100$ nodes and $E=N,2N,3N,4N$ expected edges; 2) sampling random instantiations of causal mechanisms (Linear, NN with additive/non-additive Gaussian noise); and 3) iteratively sampling observations in topological order (details in Appendix B.1). From every causal graphical model (steps 1-2), we generated two datasets, each with $1000N$ points: either all observational, or split equally among regimes (observational and perfect single-node interventions on all nodes). All models that can accommodate interventions were run on the interventional datasets, with complete knowledge of the intervention target identities. The remaining were run on the observational datasets. We generated 90 training, 5 validation, and 5 testing datasets for each combination. For testing, we also sampled out-of-distribution datasets with 1) Sigmoid and Polynomial mechanisms with Gaussian noise; and 2) Linear with additive non-Gaussian noise.
4.2 Metrics
We report standard causal discovery metrics. These include both discrete and continuous metrics, as neural networks can be notoriously uncalibrated (Guo et al., 2017), and arbitrary discretization thresholds may impact the robustness of findings (Ng et al., 2024; Schaeffer et al., 2023). For all continuous metrics, we exclude the diagonal since several baselines manually set it to zero (Brouillard et al., 2020; Lopez et al., 2022).
SHD: Structural Hamming distance is the minimum number of edge edits required to match two graphs (Tsamardinos et al., 2006) (predicted and true DAGs, or the implied CPDAGs). Discretization thresholds are as published or default to 0.5.
mAP: Mean average precision computes the area under precision-recall curve per edge and averages over the graph (predicted and true DAGs, or undirected skeletons). The random guessing baseline depends on the positive rate.
@@ 4.2 Metrics
Since OA is normalized by $\|E\|$, it is invariant to the positive rate. In contrast to orientation F1 (Geffner et al., 2024), it is also invariant to the assignment of forward/reverse edges as 1/0.
@@ 4.3 Baselines
We compare against several deep learning and classical baselines. All baselines were trained and/or run from scratch on each testing dataset using their published code and hyperparameters, except Avici (their recommended checkpoint was trained on their synthetic train and test sets after publication, Appendix B.2).
DCDI (Brouillard et al., 2020) extracts the causal graph as a parameter of a generative model. The G and Dsf variants use Gaussian or deep sigmoidal flow likelihoods, respectively. DCD-FG (Lopez et al., 2022) follows DCDI-G, but factorizes the graph into a product of two low-rank matrices for scalability. DiffAn (Sanchez et al., 2023) uses the trained model’s Hessian to obtain a topological ordering, followed by a classical pruning algorithm. AVICI (Lorch et al., 2022) uses an amortized inference approach to estimate $P(G\mid D)$ over a class of data-generating mechanisms via variational inference. VarSort (a.k.a. “sort and regress”) (Reisach et al., 2021) sorts nodes by marginal variance and sparsely regresses nodes based on their predecessors. This naive baseline is intended to reveal artifacts of synthetic data generation. FCI, GIES quantify the predictive power of FCI and GIES estimates, when run over all nodes. VarSort, Fci, and Gies were run using non-parametric bootstrapping (Friedman et al., 1999), with 100 estimates of the full graph (based on 1000 examples each), where the final prediction for each edge is its frequency of appearing as directed. Since these methods are treated as oracles, the bootstrapping strategy was selected to maximize test performance. Visualizations of the bootstrapped graph can be found in Figures 7 and 8.
@@ 5 Results
We highlight representative results in each section, with additional experiments and analyses in Appendix C.
@@
    Section 5.2 focuses on the case where we do know (or can estimate) the class of causal mechanisms or exogenous noise. We show that adapting our pretrained models with this information at zero/low cost leads to substantial improvement and exceeds the best baseline trained from scratch.
  3. 3.
    Section 5.3 analyzes Sea predictions in context of classic identifiability theory. In particular, we focus on the linear Gaussian case, and show that Sea approaches “oracle” performance (with respect to the MEC), while simply running a classic discovery algorithm cannot, on our finite datasets.
  4. 4.
    Section 5.4 contains a variety of ablation studies. In particular, Sea exhibits impressive low-data performance, requiring only 400 samples to perform well on $N=100$ datasets. We also ablate estimation hyperparameters and the contribution of marginal/global features.
Table 1: Synthetic experiments. Mean/std over 5 distinct Erdős-Rényi graphs, with metrics relative to ground truth DAG. DiffAn, VarSort, Fci, Sea(Fci) run on observational data only. Evaluation w.r.t. CPDAG and undirected skeleton in Tables 14 and 15, with additional baselines and ablations in Appendix C. $\dagger$ indicates o.o.d. setting. * indicates non-parametric bootstrapping. Runtimes with 1 CPU and 1 V100 GPU.
$N$ $E$ Model Linear NN add. Sigmoid Polynomial Overall
20 20 Dcdi-G $0.59\scriptstyle\pm.12$ $6.4\scriptstyle\pm.9$ $0.78\scriptstyle\pm.07$ $\textbf{3.0}\scriptstyle\pm.7$ $0.36\scriptstyle\pm.06$ $42.7\scriptstyle\pm.3$ $0.42\scriptstyle\pm.08$ $10.4\scriptstyle\pm.4$ 4735.7
Dcdi-Dsf 0.66±.16plus-or-minus0.66.160.$6$16 $5.2\scriptstyle\pm.3$ $0.69\scriptstyle\pm.18$ $4.2\scriptstyle\pm.5$ $0.37\scriptstyle\pm.04$ $43.2\scriptstyle\pm.4$ $0.26\scriptstyle\pm.08$ $15.7\scriptstyle\pm.2$ 3569.1
DiffAn $0.19\scriptstyle\pm.09$ $40.2\scriptstyle\pm 4.4$ $0.16\scriptstyle\pm.10$ $38.6\scriptstyle\pm 3.1$ $0.29\scriptstyle\pm.11$ $19.2\scriptstyle\pm.6$ $0.09\scriptstyle\pm.03$ $49.7\scriptstyle\pm 4.6$ 434.3
Avici $0.48\scriptstyle\pm.17$ $17.2\scriptstyle\pm.1$ $0.59\scriptstyle\pm.09$ $10.8\scriptstyle\pm.1$ $0.42\scriptstyle\pm.13$ $17.2\scriptstyle\pm.8$ $0.24\scriptstyle\pm.08$ $18.4\scriptstyle\pm.1$ 2.0
VarSort* $0.81\scriptstyle\pm.08$ $10.0\scriptstyle\pm.4$ $0.81\scriptstyle\pm.15$ $6$7 $0.50\scriptstyle\pm.13$ $16.1\scriptstyle\pm.7$ 0.33±.13plus-or-minus0.33.130.$3$13 $17.1\scriptstyle\pm.1$ 0.4
Fci* 0.66±.07plus-or-minus0.66.070.$6$07 $19.0\scriptstyle\pm.3$ $0.42\scriptstyle\pm.19$ $17.4\scriptstyle\pm.2$ $0.56\scriptstyle\pm.08$ $18.5\scriptstyle\pm.5$ $0.41\scriptstyle\pm.14$ $18.9\scriptstyle\pm.3$ 22.2
Gies* $0.84\scriptstyle\pm.08$ $7.4\scriptstyle\pm.0$ $0.79\scriptstyle\pm.07$ $9.0\scriptstyle\pm.1$ $0.71\scriptstyle\pm.10$ $12.5\scriptstyle\pm.7$ $0.62\scriptstyle\pm.09$ $13.7\scriptstyle\pm.7$ 482.1
Sea (Fci) $\textbf{0.96}\scriptstyle\pm.03$ $3.2\scriptstyle\pm.6$ $0.91\scriptstyle\pm.04$ $5.0\scriptstyle\pm.8$ $\textbf{0.85}\scriptstyle\pm.09$ $\textbf{6.7}\scriptstyle\pm.1$ $\textbf{0.69}\scriptstyle\pm.09$ $\textbf{9.8}\scriptstyle\pm.2$ 4.2
Sea (Gies) $\textbf{0.97}\scriptstyle\pm.02$ $\textbf{3.0}\scriptstyle\pm.9$ $\textbf{0.94}\scriptstyle\pm.03$ $3.4\scriptstyle\pm.4$ $0.84\scriptstyle\pm.07$ $8.1\scriptstyle\pm.8$ $\textbf{0.69}\scriptstyle\pm.12$ $10.1\scriptstyle\pm.9$ 3.0
100 400 Dcd-Fg $0.05\scriptstyle\pm.00$ $\scriptstyle 3068\pm 131$ $0.07\scriptstyle\pm.00$ $\scriptstyle 3428\pm 154$ $0.13\scriptstyle\pm.02$ $\scriptstyle 3601\pm 272$ $0.12\scriptstyle\pm.03$ $\scriptstyle 3316\pm 698$ 1838.2
Avici $0.12\scriptstyle\pm.02$ $391\scriptstyle\pm 8$ $0.17\scriptstyle\pm.01$ $407\scriptstyle\pm 19$ $0.10\scriptstyle\pm.02$ $398\scriptstyle\pm 11$ $0.03\scriptstyle\pm.00$ $402\scriptstyle\pm 19$ 9.3
VarSort* $0.80\scriptstyle\pm.02$ $224\scriptstyle\pm 10$ $0.18\scriptstyle\pm.03$ $1139\scriptstyle\pm 269$ $0.51\scriptstyle\pm.05$ $350\scriptstyle\pm 15$ $0.27\scriptstyle\pm.04$ $380\scriptstyle\pm 17$ 5.1
Sea (Fci) $0.84\scriptstyle\pm.02$ $162\scriptstyle\pm 12$ $0.04\scriptstyle\pm.00$ $403\scriptstyle\pm 16$ $0.63\scriptstyle\pm.03$ $247\scriptstyle\pm 17$ $0.34\scriptstyle\pm.04$ $\textbf{325}\scriptstyle\pm 22$ 19.2
Sea (Gies) $\textbf{0.91}\scriptstyle\pm.01$ $\textbf{116}\scriptstyle\pm 7$ $\textbf{0.27}\scriptstyle\pm.10$ $\textbf{364}\scriptstyle\pm 34$ $\textbf{0.69}\scriptstyle\pm.03$ $\textbf{218}\scriptstyle\pm 21$ $\textbf{0.38}\scriptstyle\pm.04$ $328\scriptstyle\pm 22$ 3.1
5.1 SEA generalizes to out-of-distribution, misspecified, and real datasets
@@
Table 1 summarizes our controlled experiments on synthetic data. Sea exceeds all baselines in the Linear case, which matches the models’ assumptions exactly (causal discovery algorithms and inverse covariance). In the misspecified (NN) or misspecified and out-of-distribution settings (Sigmoid, Polynomial), Sea also attains the best performance in the vast majority of cases, even though Dcdi and Avici both have access to the raw data. Furthermore, our models outperform VarSort in every single setting, while most baselines are unable to do so consistently. This indicates that our models do not simply overfit to spurious features of the synthetic data generation process.
@@
Table 3 illustrates that we exceed baselines on single cell gene expression data from CausalBench (Chevalley et al., 2025; Replogle et al., 2022). Furthermore, when we increase the subset size to $b=2000$, we achieve very high precision (0.838) over 2834 predicted edges. Sea runs within 5s on this dataset of 162k cells and $N=622$ genes, while the fastest naive baseline takes 5 minutes and the slowest deep learning baseline takes 9 hours (run in parallel on subsets of genes).
5.2 SEA adapts to new data assumptions with zero to minimal finetuning
We illustrate two strategies that allow us to use pretrained Sea models with different implicit assumptions. First, if two causal discovery algorithms share the same output format, they can be used interchangeably for marginal estimation. On observational, linear non-Gaussian data, replacing the Ges algorithm with Lingam (Shimizu et al., 2006) is beneficial without any other change (Table 5). The same improvement can be observed on Polynomial and Sigmoid non-additive data, when running Fci with a polynomial kernel conditional independence test (Kci, Zhang et al. (2011)) instead of the Fisherz test, which assumes linearity (Table 5). In principle, different algorithms might make different mistakes, so this strategy could lead to out-of-distribution inputs for the pretrained aggregator. While comparing across algorithms is not a primary focus of this work and requires further theoretical study, we notice similar performance for alternate estimation algorithms with linear Gaussian assumptions (Table 5.2), regardless of discovery strategy (GRaSP (Lam et al., 2022)). The gap is larger for LiNGAM, which assumes non-Gaussianity (Table 25).
@@ 5.2 SEA adapts to new data assumptions with zero to minimal finetuning
Table 6: Sea respects identifiability theory. Observational setting, standardized (-std) $N=10,E=10$ linear Gaussian test datasets with $>1$ graph in Markov equivalence class (MEC). Top: oracle performance based on true MEC (see left). Bottom: trained Sea approaches oracle performance, while FCI is very noisy.
Model mAP($\uparrow$) AUC($\uparrow$) SHD($\downarrow$) OA($\uparrow$)
metric over mean MEC 0.88±.10plus-or-minus0.88.100.$8$10 $0.98\pm\scriptstyle.03$ $2.0\pm\scriptstyle 1.0$ $0.74\pm\scriptstyle.22$
mean metric over MEC $0.74\pm\scriptstyle.21$ $0.91\pm\scriptstyle.07$ $1.2\pm\scriptstyle.69$ $0.84\pm\scriptstyle.13$
Sea(Fci)-std $0.83\pm\scriptstyle.16$ $0.97\pm\scriptstyle.04$ $3.3\pm\scriptstyle 2.3$ $0.69\pm\scriptstyle.21$
Sea(Fci)+Corr-std $0.84\pm\scriptstyle.14$ $0.96\pm\scriptstyle.03$ $5.0\pm\scriptstyle 4.5$ $0.85\pm\scriptstyle.14$
Fci-std $0.49\pm\scriptstyle.28$ $0.75\pm\scriptstyle.16$ $9.3\pm\scriptstyle 2.8$ $0.49\pm\scriptstyle.29$
Table 7: Sea is generally insensitive to swapping estimation other algorithms with linear Gaussian assumptions, at inference time. Results on $N=10$ observational setting. FCI cannot be used with Sea(g) since FCI outputs a PAG, not a CPDAG.
Inference estimator Sea (Fci) Sea (Gies)
Lin. NN Sig. Poly. Lin. NN Sig. Poly.
FCI $0.98$ $0.88$ $0.83$ $0.62$
PC $0.93$ $0.85$ $0.86$ $0.64$ $0.96$ $0.89$ $0.82$ $0.58$
GES $0.94$ $0.85$ $0.80$ $0.60$ $0.95$ $0.88$ $0.81$ $0.57$
GRaSP $0.93$ $0.85$ $0.80$ $0.61$ $0.95$ $0.88$ $0.81$ $0.57$
Figure 3: Few-shot learning behavior emerges as training set increases. “Tiny” Sea trained on 1/4 of the data is comparable to the full model on $N=10$ datasets when given $T=50$ batches, but is less robust with only $T=10$.
Figure 4: Ablations with Sea (Gies) for estimation parameters on $N=100,E=100$. Error bars indicate 95% confidence interval across the 5 i.i.d. datasets of each setting. All parameters are set to the defaults (Section 3.3) unless otherwise noted. (A) Dashed: inverse covariance at $M=500$. (C) Variance is unusually high for Sigmoid $b=300$ until $T=100$, indicating that larger batches result in more stable results.
Table 8: Ablating marginal and global features on Sea (Gies). Top: We set marginal and global representations to 0 (lack of $E^{\prime}$/$\rho$ is out-of-distribution) and observe that the pretrained model is more robust to removing $E^{\prime}$, perhaps since we sample varying $T$ during training. Bottom: Re-train Sea (Gies) on $N=10$ datasets, with and without global features (lack of $\rho$ is in-distribution). We observe that global features are “worth” $T\approx 40$ estimates of $k=5$ variables each.
Model Linear NN add. NN. Sigmoid Polynomial
Sea (Gies) 0.99±.01plus-or-minus0.99.010.$9$01 $1.2\scriptstyle\pm.7$ $0.94\scriptstyle\pm.06$ $2.6\scriptstyle\pm.8$ $0.91\scriptstyle\pm.07$ $3.2\scriptstyle\pm.3$ $0.85\scriptstyle\pm.12$ $4.0\scriptstyle\pm.5$ $0.70\scriptstyle\pm.11$ $5.8\scriptstyle\pm.6$
$h^{\rho}\leftarrow 0$ $0.30\scriptstyle\pm.17$ $29.2\scriptstyle\pm.4$ $0.27\scriptstyle\pm.18$ $29.4\scriptstyle\pm.8$ $0.19\scriptstyle\pm.09$ $29.0\scriptstyle\pm.0$ $0.35\scriptstyle\pm.17$ $27.4\scriptstyle\pm.4$ $0.31\scriptstyle\pm.15$ $27.1\scriptstyle\pm.9$
$h^{E}\leftarrow 0$ $0.85\scriptstyle\pm.09$ $6.4\scriptstyle\pm.7$ $0.82\scriptstyle\pm.11$ $10.2\scriptstyle\pm.9$ $0.78\scriptstyle\pm.07$ $13.2\scriptstyle\pm.2$ $0.63\scriptstyle\pm.21$ $10.2\scriptstyle\pm.8$ 0.55±.19plus-or-minus0.55.190.$5$19 $13.4\scriptstyle\pm.6$
$T=2$ $0.20\scriptstyle\pm.04$ $31.2\scriptstyle\pm.5$ $0.25\scriptstyle\pm.06$ $29.2\scriptstyle\pm.4$ 0.33±.10plus-or-minus0.33.100.$3$10 $27.8\scriptstyle\pm.1$ $0.19\scriptstyle\pm.07$ $30.2\scriptstyle\pm.9$ $0.24\scriptstyle\pm.09$ $28.1\scriptstyle\pm.9$
$T=10$ $0.62\scriptstyle\pm.16$ $8.0\scriptstyle\pm.8$ $0.69\scriptstyle\pm.11$ $9.6\scriptstyle\pm.6$ 0.66±.13plus-or-minus0.66.130.$6$13 $11.2\scriptstyle\pm.9$ $0.62\scriptstyle\pm.20$ $9.2\scriptstyle\pm.6$ $0.50\scriptstyle\pm.22$ $8.9\scriptstyle\pm.2$
$T=50$, no $\rho$ $0.63\scriptstyle\pm.13$ $6.8\scriptstyle\pm.1$ $0.53\scriptstyle\pm.07$ $6.2\scriptstyle\pm.9$ $0.68\scriptstyle\pm.20$ $6.0\scriptstyle\pm.1$ $0.58\scriptstyle\pm.15$ $7.1\scriptstyle\pm.1$ $0.50\scriptstyle\pm.14$ $7.1\scriptstyle\pm.5$
Another strategy is to “finetune” the aggregator, either fully or using low-cost methods like LoRA (Hu et al., 2022). Specifically, we keep the same training set and classification objective, while changing the input’s featurization, e.g. a different global statistic. Here, we show that finetuning our models for distance correlation (Dcor) is beneficial in both Tables 5 and 5, and the combination of strategies results in the highest performance overall, surpassing the best baseline trained from scratch (Dcdi-Dsf).